OpenStudy (anonymous):

Factor completely: 2x3 + 4x2 + 6x + 12

OpenStudy (anonymous):

2[(x^3 + 3x) + (2x^2 + 6)] = 2[x*(x^2 + 3) + 2(x^2 + 3) ]= = 2(x^2 + 3)(x + 2)